# Solving Robotics Problems in Zero-Shot with Vision-Language Models

### v3

@@ Abstract
We introduce Wonderful Team, a multi-agent Vision Large Language Model (VLLM) framework designed to solve robotics problems in a zero-shot regime. In our context, zero-shot means that for a novel environment, we provide a VLLM with an image of the robot’s surroundings and a task description, and the VLLM outputs the sequence of actions necessary for the robot to complete the task. Unlike prior work that requires fine-tuning parts of the pipeline – such as adjusting an LLM on robot-specific data or training separate vision encoders – our approach demonstrates that with careful engineering, a single off-the-shelf VLLM can autonomously handle all aspects of a robotics task, from high-level planning to low-level location extraction and action execution. Crucially, compared to using GPT-4o alone, Wonderful Team is self-corrective and capable of iteratively fixing its own mistakes, enabling it to solve challenging long-horizon tasks. We validate our framework through extensive experiments, both in simulated environments using VIMABench and in real-world settings. Our system showcases the ability to handle diverse tasks such as manipulation, goal-reaching, and visual reasoning—all in a zero-shot manner. These results underscore a key point: vision-language models have progressed rapidly in the past year and should be strongly considered as a backbone for many robotics problems moving forward.
@@ 1 Introduction
Advancements in Large Language Models (LLMs) and Vision-Language Models (VLLMs) have brought us closer to enabling robots to perform complex tasks based solely on natural language instructions, without prior training. By integrating vision and language, VLLMs allow robots to intuitively understand their environments, leveraging real-world priors from large-scale data. However, developing a general-purpose robotic system capable of executing complex tasks in dynamic settings remains challenging. Such systems need to perceive surroundings, utilize appropriate skills, and achieve long-horizon subgoals. This raises a crucial question: Can these models be adapted to solve robotic tasks in unstructured environments without any training?
Current approaches in language-conditioned robotics often separate the problem into high-level planning and low-level perception-action execution, utilizing distinct modules for each component. While this separation can facilitate zero-shot operation, it may hinder seamless integration between perception and action, especially when modules are disconnected.
These limitations hinder the seamless integration of perception and action, as vision models like CLIP, which primarily offer class-level predictions, lack the deep environmental understanding needed for complex, context-specific tasks. Similarly, while LangSAM can segment objects based on language prompts, it struggles with precise object identification in complex scenes or when handling abstract instructions that require deeper comprehension. As a result, these models perform well with easily identifiable objects but face challenges when handling abstract or environment-specific tasks, which significantly limits their ability to help LLMs accurately ground environmental context and generate actionable outputs. The separation of planning and perception hinders the seamless integration of perception and action in decision-making. However, with the multimodal capabilities of modern VLLMs, this division may no longer be necessary. In this paper, we introduce Wonderful Team: a zero-shot, single-model, multi-agent system that unifies planning and perception within a VLLM framework using interconnected specialized agents. This integrated approach enables end-to-end reasoning and execution without relying on external modules or fine-tuning, effectively addressing the limitations of previous modular methods.
Our key contributions include:
@@
  • Zero-Shot Coordinate-Level Control in Complex Robotics Tasks: Our system operates without any prior training, fine-tuning, or environment-specific prompts, successfully handling diverse tasks in both simulated and real-world environments. It delivers precise, coordinate-level control for robotic execution, outperforming methods that rely on coarse object-level or sub-task-level instructions.
@@
  • Introducing a Multi-Agent VLLM Framework to Overcome Previous Limitations: We have developed a novel multi-agent structure within a single VLLM, where specialized agents collaboratively handle various aspects of robotic tasks, from high-level planning to low-level execution. By integrating perception and action, and employing a divide-and-conquer approach with reflection capabilities, we address the shortcomings of previous models, including issues with context-aware object identification, precise localization, and handling multiple instances of the same object.
@@
  • Empirical Validation through Extensive Experiments and Ablation Studies: We validate our framework with comprehensive experiments in both simulation (VIMABench) and real-world settings. Our results show significant performance improvements over existing methods, including those that require training. We also conduct thorough ablation studies to examine the effects of different agents and configurations, highlighting the critical role of the multi-agent system in achieving optimal performance.
@@
Demonstration videos of the robotic policies in action, along with the code, can be accessed on our project website.
@@ 2 Motivating Examples
Developing robotic systems that can understand and execute complex tasks in unstructured environments remains a significant challenge. Existing frameworks often employ a Large Language Model (LLM) as a text planner combined with a separate vision model (e.g., CLIP, OWL-ViT, LangSAM) to perceive the environment. While this modular approach seems logical, it faces critical limitations when applied to intricate, context-dependent tasks.
@@ 2.1 Can an LLM as a Planner with a Separate Vision Model Find Objects?
Not Always. There are limitations at both the planning and perception levels:
@@
Not within the current framework.
Even with enhanced reasoning and replanning, we are unable to fully address LangSAM’s limitations because the LLM lacks the capability to detect, notice, or correct errors originating from the seperate vision model.
@@ 2.2 Could simply replacing LangSAM with a VLLM resolve these issues?
Partially; a VLLM may improve context comprehension, but it fails to match the precision that LangSAM already provides.
To provide a clearer context for spatial reasoning, we first introduce pixel coordinates as a reference framework (see Figure 2). Without this grid overlay, even humans might struggle to describe relative locations accurately in a complex scene.
Figure 2: An example scenario with overlaid pixel coordinates.
However, there are still notable challenges with this framework:
@@
2. Difficulty with Complex Instructions: Tasks that require understanding spatial relationships or handling multiple objects can overwhelm the reasoning capabilities. Observation 1: VLLMs Can Recognize and Diagnose Their Own Errors
VLLMs have the ability to detect mistakes in their outputs and adjust them upon review. For example, when asked to locate a cluster of grapes, the model may initially provide an imprecise answer, but can correct it when prompted to reassess (see Figure 3). Table 7 shows GPT-4o’s 97% success in classifying bounding boxes, highlighting its self-assessment abilities. This suggests VLLMs can iteratively refine outputs, even from initially imprecise coordinates.
@@
Observation 2: VLLMs Can Self-Correct Through Reflection
@@
VLLMs can iteratively refine their outputs based on feedback, a process known as reflection. Over several iterations, they improve their estimation of an object’s position, moving closer to the correct target (see Figure 4).
While using a VLLM alone naively is insufficient, these observations reveal the potential to address its limitations by leveraging its self-correction capabilities in a structured way.
@@ 3 Wonderful Team
Building on these insights, we propose a novel pipeline for robotics that leverages specialized agents, each responsible for a distinct part of the reasoning process within a structured framework. By combining the strengths of Vision-Language Models (VLLMs) and breaking down complex tasks into manageable components, each agent can focus on a specific role, resulting in more precise and reliable robotic control. As illustrated in Figure 5, our multi-agent framework defines the distinct roles of each agent, the flow of information from high-level tasks to low-level actions, and their collaborative efforts in executing tasks effectively.
Each agent in our system is designed to address specific challenges in robotic tasks. For example, in Figure 5(b), when the robot is instructed to “put the banana into the box,” the initial plan generated by the Supervisor agent often overlooks obstacles like the box’s lid. This is where the Verification agent plays a critical role. Its reflection process involves reviewing the subgoal plan, checking for potential issues such as physical constraints or incomplete steps, and cross-referencing this plan with the current state of the environment. If an issue, like the lid blocking access to the box, is detected, the Verification agent raises this concern to the Supervisor. This early feedback allows the system to refine the plan before executing any action. Unlike the replanning process, which occurs at the end of the pipeline if a task fails, the Verification agent catches errors early to prevent failures and avoid costly adjustments later. This proactive approach enhances the robustness and adaptability of the robotic control.
@@
The Grounding team then takes over to refine the coordinates for each target, ensuring precise and collision-free movements. The Mover and Checker agents collaborate through an iterative process of adjusting positional groundings. Figure 4 provides an example of the Grounding team in action. The separation of tasks into a multi-agent system proves advantageous, as it allows each agent to focus on its distinct responsibilities with varying levels of access to critical information. For a detailed discussion on the benefits of this multi-agent approach, refer to Appendix E.4.
Are all parts of the Wonderful Team necessary? Ablation studies reveal that all components of the Wonderful Team are essential. Removing memory agents leads to failures, such as mistaking irrelevant objects for targets, while omitting grounding members results in inaccurate coordinates. A supervisor-only setup works for simple tasks but fails with complex ones, lacking precision and corrective processes. Appendix C provides detailed analysis, and Table 4 in the appendix shows the impact on success rates when specific agents are removed.
@@ 4 Related Work
Language Models Empowering Robotics: Prior work has leveraged natural language to enhance robotic learning and adaptation. Early approaches equipped agents with learned language embeddings, requiring large amounts of training data (Bing et al., 2023; Jiang et al., 2023). Others focused on connecting language instructions with low-level action primitives to solve long-horizon tasks (Hu et al., 2023; Huang et al., 2022b; Liang et al., 2023). While effective in specific contexts, these methods often struggle to generalize to new tasks without retraining. Foundation models like RT-1 (Brohan et al., 2022) and RT-2 (Brohan et al., 2023) have advanced versatile robotic systems, but they still require significant training to achieve robust performance across diverse tasks.
@@
Vision-Language Models for Localization: PIVOT (Nasiriany et al., 2024) addresses enabling VLMs to localize actionable points without fine-tuning on task-specific data. Their approach centers on localization through visual question answering, with minimal focus on planning—similar to the role of our Grounding Team. Unlike our method, which integrates both localization and planning within a multi-agent framework, PIVOT primarily addresses localization without managing complex, long-horizon tasks. In PIVOT, a single agent iteratively selects action points, whereas our approach employs multiple agents with distinct roles for refining and verifying actions. A detailed comparison is provided in Appendix E.2.
Language Models as Zero-Shot Trajectory Generators: Kwon et al. (2024) propose using language models as zero-shot trajectory generators. Their approach uses a predefined object detection model (LangSAM) to extract object information, which is then used by the LLM to plan. Specifically, the LLM generates Python scripts to create a trajectory for execution. Unlike our method, which uses a VLLM to integrate perception and action without external modules, their approach relies on separate perception models and code generation for trajectory planning. Further comparison is available in Appendix E.3.
Natural Language as Policies: Concurrent with our work, Natural Language as Policies (NLaP) (Mikami et al., 2024) developed a few-shot, end-to-end model for coordinate-level action prediction. Their approach involves providing a one-shot example, either from the same task or a closely related one, rather than adopting a zero-shot paradigm. Unlike our method, which integrates both grounding and planning within a multi-agent framework, NLaP focuses less on grounding and directly uses system information from the environment, bypassing the need to extract coordinates from images using VLMs. NLaP serves as one of the baselines in our experiments, and a detailed comparison is presented in Appendix E.1.
@@ 5.4 Results and Discussion
(a) Comparison with baseline methods. Grey boxes indicate reduced complexity due to the framework’s nature, which should be considered when interpreting results.
@@
As shown in Figure 9(c), Wonderful Team outperforms baselines across all VIMABench tasks. The Grounding Team and multi-agent structure are crucial; removing or replacing them significantly reduces performance. Methods like Trajectory Generator and our ablation with a separate VLM struggle to detect uncommon objects and lack nuanced reasoning for detection and manipulation. Even with perfect localization (as in NLaP), complex long-horizon planning remains challenging without the multi-agent structure, leading to misinterpretations and errors (Appendix E.1). Ablation studies (Appendix C) confirm the importance of each component in Wonderful Team.

### v5

# Wonderful Team: Zero-Shot Physical Task Planning with Visual LLMs

@@ Abstract
We introduce Wonderful Team, a multi-agent Vision Large Language Model (VLLM) framework for executing high level robotic planning in a zero-shot regime. In our context, zero-shot high-level planning means that for a novel environment, we provide a VLLM with an image of the robot’s surroundings and a task description, and the VLLM outputs the sequence of actions necessary for the robot to complete the task. Unlike previous methods for high-level visual planning for robotic manipulation, our method uses VLLMs for the entire planning process, enabling a more tightly integrated loop between perception, control, and planning. As a result, Wonderful Team’s performance on a real-world semantic and physical planning tasks often exceeds methods that rely on separate vision systems. For example, we see an average 40% success-rate improvement on VimaBench over prior methods such as NLaP, an average 30% improvement over Trajectory Generators on tasks from the Trajectory Generator paper including drawing and wiping a plate, and an average 70% improvement over Trajectory Generators on a new set of semantic reasoning tasks including environment re-arrangement with implicit linguistic constraints. We hope these results highlight the rapid improvements of VLLMs in the past year, and motivate the community to consider VLLMs as an option for some high-level robotic planning problems in the future.
@@ 1 Introduction
High-level robotic planning is a difficult task to pin down. It requires an extensive knowledge of the physical world. If planning is to be done via language instructions, it also requires a semantic understanding of the language itself, and how the objects in the environment relate to sometimes vague linguistic instructions. Consider, for example, the task “Pick up the object to the right of the grapes and put it in the box.” To solve this task, we need to identify ‘the object to the right of the grapes,’ which is a non-trivial visual relationship inference problem. We also need to consider the implicit constraints of the environment. For example, if the box has a lid and is currently closed, we need to infer that the lid must first be removed before proceeding with the task.
Figure 1: Comparison of high-level planning approaches for a color-matching fruit placement task. MOKA’s planned trajectory (b) shows several limitations: its unverified high-level plan generates suboptimal subgoals due to the lack of self correction (step 2 treating grape as green) and shows misalignment between VLLM-generated targets and visual grounding. For instance, at step 3, it identifies the orange fruit to be the ‘orange area’, and at step 4, it fails to properly pick the green apple due to not specifying the color. We discuss more on comparision with MOKA in Appendix F.4.
Recently, Large Language Models (LLMs) have emerged as an enticing pathway for achieving high-level robotic planning. LLMs already have a built-in world model, including priors over object names and relative relationships. They also have the ability to understand and parse semantic meaning from language instructions. Several promising pipelines exist for converting language commands to high-level robotics plans Kwon et al. (2024). Typically, if vision-based planning is required, these pipelines make use of an off-the-shelf vision encoder such as LangSAM or OWL-ViT that finds objects in an environment and delivers their coordinates in pixel space. While this is a great idea in theory, we find that, in practice, such reliance on an external vision system often results in pipelines that are brittle and unable to recover from mistakes originating in the visual system.
In this paper, we consider the problem of using Visual Large Language Models (VLLMs) to generate action-level robotics plans directly from environmental images. Compared to prior works, we find having VLLMs as the backbone of both planning and grounding in the pipeline to have several advantages. First, VLLMs combine the semantic reasoning power of language models with robust visual understanding, bridging the traditional gap between high-level planning and visual grounding. Thus, they do not struggle with commands such as “pick the orange and place it in the orange area.” The model can read the environment image and understand implicitly to treat the orange square as the orange area instead of the orange fruit, although both can be considered “orange area” without context. More importantly, VLLMs enable self-correction at both the planning and perception levels. For planning, if the initial high-level plan overlooks environmental constraints, VLLMs can identify these errors and suggest corrective actions. For perception, while traditional vision models fail permanently when objects cannot be found, our VLLM-based perception module can refine its search strategy through reasoning about alternative viewpoints and object appearances. This ability to self-correct improves success rates on real-world spatial and semantic planning tasks by approximately 80% versus a baseline GPT-4o model with no self-correction, with many tasks going from a 0% success rate to a 100% success rate after self-correction is introduced.
Our contributions can be summarized as follows
@@
  • Zero-Shot Coordinate-Level Control in Complex Robotics Tasks: Our system operates without any prior training, fine-tuning, or environment-specific prompts, successfully handling diverse tasks in both simulated and real-world environments.
@@
  • Developed a Multi-Agent VLLM Framework to Overcome Previous Limitations in High-Level Robotic Planning from Language Instructions: We have developed a novel multi-agent structure within a single VLLM, where specialized agents collaboratively handle various aspects of high level robotic planning, from sequential position generation to location estimation. By integrating perception and action, and employing a divide-and-conquer approach with reflection capabilities, we address the shortcomings of previous models, including issues with context-aware object identification, precise localization, and handling multiple instances of the same object.
@@
  • Empirical Validation through Extensive Experiments and Ablation Studies: We validate our framework with comprehensive experiments in both simulation (VIMABench) and real-world settings, including spatial planning tasks from Trajectory Generators and a new set of semantic reasoning real-world robotics tasks. Our results show significant performance improvements over existing methods on problems related to physical task planning. Interestingly, these improvements are also present when comparing against planning methods that require training. We also conduct thorough ablation studies to examine the effects of different agents and configurations, highlighting the critical role of the multi-agent system in achieving optimal performance.
@@
Demonstration videos of the robotic policies in action, along with the code and prompts, can be accessed on our project website.
@@ 2 Motivating Examples
Developing robotic systems that can understand and execute planning over complex manipulation tasks remains a significant challenge, particularly when dealing with uncommon objects and abstract visual attributes. Existing frameworks often employ a Large Language Model (LLM) as a text planner combined with a separate vision model (e.g., CLIP, OWL-ViT, LangSAM, GroundingDINO) to perceive the environment. While this modular approach seems logical, it faces critical limitations when handling less common objects, interpreting abstract visual relationships, or reasoning about semantic properties - challenges where our integrated approach shows significant improvements over prior methods.
@@ 2.1 Can an LLM as a Planner with a Separate Vision Model Find Objects?
Not Always. There are limitations at both the planning and perception levels, which are both intimately ralted.
@@
Not within the current framework. Even with enhanced reasoning and replanning, we are unable to fully address LangSAM’s limitations because the LLM lacks the capability to detect, notice, or correct errors originating from the seperate vision model.
@@ 2.2 Could simply replacing LangSAM with a VLLM resolve these issues?
The answer to this question is ‘no.’ While replacing LangSAM with a VLLM may improve context comprehension, the resulting system fails to match the precision that LangSAM already provides. And thus the substitution is largely a wash. There are a few reasons for this.
@@
2. Difficulty with Complex Instructions: Tasks that require understanding spatial relationships or handling multiple objects can overwhelm the reasoning capabilities.
However, while experimenting with using VLLMs to generate high-level plans, we had an interesting observation: VLLMs often know they’re wrong. And they can often diagnose their own errors.
For example, when asked to locate a cluster of grapes, the model may initially provide an imprecise answer, but can correct it when prompted to reassess (see Figure 3). Table 14 shows GPT-4o’s 97% success in classifying bounding boxes, highlighting its self-assessment abilities. This suggests VLLMs can iteratively refine outputs, even from initially imprecise coordinates.
@@
Even more interesting than this first observation, we also noticed that VLLMs often know how to iteratively self-correct errors in their perception. Over several iterations, they can improve their estimation of an object’s position, moving closer to the correct target (see Figure 4). This process of an LLM iteratively refine its outputs based on feedback is known as reflection.
@@
Thus, we arrive at our main thesis. While naively using a VLLM for high-level robotic planning is often insufficient, the effectiveness of VLLMs for high-level robotic planning can be improved substantially if we leverage their self-correction capabilities in a structured way.
@@ 3 Wonderful Team
Building on insights from the previous section, we propose a novel pipeline for high-level physical task planning in robotics that leverages specialized agents, each responsible for a distinct part of the reasoning process within a structured framework. By combining the strengths of Vision-Language Models (VLLMs) and breaking down complex tasks into manageable components, each agent can focus on a specific role, resulting in more precise and reliable high-level planning. As illustrated in Figure 6, our multi-agent framework defines the distinct roles of each agent, the flow of information from high-level tasks to low-level actions, and their collaborative efforts in executing tasks effectively.
We discuss the role of each team member and the flow of information between them below. Simplified prompts and more details on the implementation and workflow for each team member can be found in Appendix B. Full prompts are available in the project’s codebase.
Figure 5: Overview of the major components of the Wonderful Team. Each part of the pipeline receives different levels of input, with a unique scope and specialization within the project. The agents collaboratively handle tasks ranging from high-level planning and logical verification to precise spatial reasoning and memory management, ensuring robust and efficient execution.
Each agent in our system is designed to address specific challenges in high level planning for physical manipulation tasks. For example, in Figure 6(b), when the robot is instructed to “put the banana into the box,” the initial plan generated by the Supervisor agent often overlooks obstacles like the box’s lid. This is where the Verification agent plays a critical role. Its reflection process involves reviewing the subgoal plan, checking for potential issues such as physical constraints or incomplete steps, and cross-referencing this plan with the current state of the environment. If an issue, like the lid blocking access to the box, is detected, the Verification agent raises this concern to the Supervisor. This early feedback allows the system to refine the plan before executing any action.
The Grounding team then takes over to refine the coordinates for each target, ensuring precise and collision-free movements. The Mover and Checker agents collaborate through an iterative process of adjusting positional groundings. Figure 4 provides an example of the Grounding team in action. The separation of tasks into a multi-agent system proves advantageous, as it allows each agent to focus on its distinct responsibilities with varying levels of access to critical information. For a detailed discussion on the benefits of this multi-agent approach, refer to Appendix F.3.3.
Are all parts of the Wonderful Team necessary? Ablation studies reveal that all components of the Wonderful Team are essential. Removing memory agents leads to failures, such as mistaking irrelevant objects for targets, while omitting grounding members results in inaccurate coordinates. A supervisor-only setup works for simple tasks but fails with complex ones, lacking precision and corrective processes. Appendix D provides detailed analysis, and Table 11 in the appendix shows the impact on success rates when specific agents are removed. Appendix B shows the information flow from one agent to the next within the team.
@@ 4 Related Work
Language Models Empowering Robotics: Prior work has leveraged natural language to enhance robotic learning and adaptation. Early approaches equipped agents with learned language embeddings, requiring large amounts of training data (Bing et al., 2023; Jiang et al., 2023). Others focused on connecting language instructions with low-level action primitives to solve long-horizon tasks (Hu et al., 2023; Huang et al., 2022b; Liang et al., 2023). While effective in specific contexts, these methods often struggle to generalize to new tasks without retraining. Foundation VLA models like RT-1 (Brohan et al., 2022), RT-2 (Brohan et al., 2023), and OpenVLA (Kim et al., 2024) have advanced versatile robotic systems, but they still require significant training to achieve robust performance across diverse tasks. Additionally, from a system design perspective, directly mapping vision to action embeddings in these approaches is less intuitive compared to breaking down the problem into interpretable subtasks or subgoals, making it harder to intuitively reason about and improve the system’s behavior.
@@
Vision-Language Models for Localization: PIVOT (Nasiriany et al., 2024) addresses enabling VLMs to localize actionable points without fine-tuning on task-specific data. Their approach centers on localization through visual question answering, with minimal focus on planning—similar to the role of our Grounding Team. Unlike our method, which integrates both localization and planning within a multi-agent framework, PIVOT primarily addresses localization without managing complex, long-horizon tasks. In PIVOT, a single agent iteratively selects action points, whereas our approach employs multiple agents with distinct roles for refining and verifying actions. A detailed comparison is provided in Appendix F.2.
Language Models as Zero-Shot Trajectory Generators and MOKA: Recent works by Kwon et al. (2024) and Fang et al. (2024) demonstrate the use of pre-trained large language models (LLMs) for generating affordance-based, point-level subgoal trajectories. Both approaches rely on dedicated object detection systems that use a combination of GroundingDINO and SAM to extract object information, which informs LLM-based planning. These methods employ LLMs to generate executable Python scripts that define the robot’s trajectory. In contrast, our approach fundamentally differs by leveraging the same vision large language model for perception to seamlessly integrate planning and grounding without depending on external modules. We provide detailed comparisons in Appendices F.3 and F.4.
Natural Language as Policies: Concurrent with our work, Natural Language as Policies (NLaP) (Mikami et al., 2024) developed a few-shot, end-to-end model for coordinate-level action prediction. Their approach involves providing a one-shot example, either from the same task or a closely related one, rather than adopting a zero-shot paradigm. Unlike our method, which integrates both grounding and planning within a multi-agent framework, NLaP focuses less on grounding and directly uses system information from the environment, bypassing the need to extract coordinates from images using VLMs. NLaP serves as one of the baselines in our experiments, and a detailed comparison is presented in Appendix F.1.
@@ 5.4 Results and Discussion
(a) Comparison with Baseline Methods. Grey boxes highlight areas of reduced complexity attributable to the framework’s inherent design. This reduction should be factored into the interpretation of results. It is important to note that LangSAM is a variant that integrates GroundingDino and SAM, so both the Trajectory Generator and MOKA leverage the same base models for perception tasks.
@@
As shown in Figure 10(c), Wonderful Team outperforms baselines across all VIMABench tasks. The Grounding Team and multi-agent structure are crucial; removing or replacing them significantly reduces performance. Methods like Trajectory Generator and our ablation with a separate VLM struggle to detect uncommon objects and lack nuanced reasoning for detection and manipulation. Even with perfect localization (as in NLaP), complex long-horizon planning remains challenging without the multi-agent structure, leading to misinterpretations and errors (Appendix F.1). Ablation studies (Appendix D) confirm the importance of each component in Wonderful Team.